OpenStudy (anonymous):

The height of a trapezoid is 13 inches, one of the bases measures 22 inches. If the area of the trapezoid is 318.5 square inches, find the length of the other base.

OpenStudy (anonymous):

A. 2.5 inches B. 71 inches C. 34.25 inches D. 27 inches

OpenStudy (anonymous):

@Callisto

OpenStudy (callisto):



OpenStudy (anonymous):

44*13?

OpenStudy (callisto):

Let x be another base area = (x + 22) x height /2 You have area and height. Can you plug in the numbers there and solve for x?

OpenStudy (anonymous):

area=318.5 318.5=(x+22)x heoght 2 do i go from there?

OpenStudy (anonymous):

ahhh i can awnser its to hard >.<

OpenStudy (callisto):

What is the height?

OpenStudy (anonymous):

13inc

OpenStudy (callisto):

So, you get 318.5 = (x+22) x13 Do you know how to solve it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

32.5

OpenStudy (anonymous):

thx a lot

OpenStudy (anonymous):

wait dont i do (13x+286)

OpenStudy (callisto):

318.5 = (x+22) x13 You can first divide both sides by 13, then subtract 22 from both sides.

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

2.5

OpenStudy (anonymous):

mind helping with 2 more?

OpenStudy (anonymous):

Quadrilateral MATH is a square. If MA = 2x – 5 and AT = x + 10, find the perimeter of the square

OpenStudy (anonymous):

can u help?

OpenStudy (anonymous):

Well you set 2x-5=x+10